# Game-Theoretic Joint Incentive and Cut Layer Selection Mechanism in Split Federated Learning

### v1

@@ Abstract
Split Federated Learning, Game Theoryformatting, style, styling, insert
@@ I Introduction
  • FL: FL allows a server model to be trained in parallel across many clients, but each client is required to run an entire model, resulting in a significant training burden. This leads to high battery power consumption on clients, significant client-side computation requirements, and limited model privacy.
@@
  • Our findings show that the Nash and Stackelberg strategies are desirable operating points for all participants (i.e., the SFL model owner and clients) in competitive situations. From our classification experiments using the CIFAR-10 dataset, we highlight the importance of incentive mechanisms in accelerating SFL and enhancing model accuracy. In our experiment, accuracy varies from 56% to 87% depending on the incentives provided.
The remainder of this paper is organized as follows. Section II presents the system model and formulates the utility functions for both the SFL model owner and the clients. In Section III, we describe the competitive game between the SFL model owner and the clients, rigorously formulating the problem as a Stackelberg game. We then characterize the closed-form solutions for the Nash Equilibrium of the clients and provide the algorithm to obtain the Stackelberg Equilibrium. Section IV discusses the impact of cut layer selection and differential privacy on the privacy level, offering guidelines for setting the minimum cut layer requirements during the SFL process. Section V reviews related work on cut layer selection and resource management for SFL. Finally, Section VI concludes the paper and suggests directions for future research.
@@
II System Model
@@
The SFL framework consists of two servers (that can be either physical or logical servers): (i) a federated learning server, referred to as the fed server, and (ii) a main server, along with multiple clients participating as part of the SFL process. In this setup, we define a set of clients as $\mathcal{N}={1,2,\cdots,N}$, where each client is represented as $n\in\mathcal{N}$ and has its own local dataset $\mathcal{D}_{n}$. The size of each dataset $\mathcal{D}_{n}$ is indicated by the number of data items, denoted as $D_{n}=|\mathcal{D}_{n}|$, with each data item comprising both features and labels. The entire model, denoted as $w$, is composed of $L$ layers. Specifically, the model $w$ is divided into two sub-models at the $L_{c}$-th layer, referred to as the cut layer. These sub-models are: (i) the client-side model $w_{c}$, including layers from the first layer up to layer $L_{c}$, and (ii) the server-side model $w_{s}$, encompassing layers from layer $L_{c}+1$ to layer $L$. Thus, the full model can be represented as $w=[w_{c};w_{s}]$. As shown in Fig. 1, each client $n$ in the set $\mathcal{N}$ maintains its client-side model, denoted as $w_{c,n}$. These client-side models are trained in parallel using the clients’ respective local datasets, through interactions with the server-side model $w_{s}$ hosted on the main server. To synchronize the client-side models, clients participate in federated learning, coordinated by the fed server, to construct a global client-side model $w_{c}$ in a collaborative manner [10].
@@
Finally, the entire SFL process is carried out over $T$ rounds. At the beginning of each round $t$, clients are provided with the latest global client-side model $w_{c}$ from the fed server. Following this, both the clients and the main server engage in training over $E$ epochs, updating their respective client-side models $w_{c,n}$ for each client $n\in\mathcal{N}$, and the server-side model $w_{s}$. Then, each client $n$ randomly samples a mini-batch instance $\zeta_{n}$ from its dataset $\mathcal{D}_{n}$. Note that during each epoch, clients randomly partition their datasets into $M$ mini-batches. Stochastic gradient descent (SGD) is applied to each mini-batch, allowing clients to update their client-side models $w_{c,n}$ and the main server to update the server-side model $w_{s}$ in a coordinated fashion. Once the $E$ epochs are concluded, the fed server collects the updated client-side models $w_{c,n}$ from all clients and combines them to form an updated global client-side model $w_{c}$ [^1] With mini-batch instance $\zeta_{n}$, $l_{n}(w;\zeta_{n})$ denotes the loss of the model $w$ over client $n$’s mini-batch instance $\zeta_{n}$, which can be expressed as follows: $l_{n}(w;\zeta_{n})=h(w_{s};u(w_{c,n};\zeta_{n})).$ (1) Here, $u$ is a function on the client side that transforms the input data from the sampled mini-batch instance $\zeta_{n}$ into the activation space, producing the output of the client-side model $w_{c,n}$ (i.e., smashed data). Meanwhile, $h$ is a function on the server side that takes the activation space as input and produces a single scalar value representing the loss..
II-A Utility Function of SFL Model Owner
@@
$U_{MO}(R,L_{c},\textbf{d})=\tau_{1}\ln(1+\frac{\sum_{n=1}^{N}d_{n}}{d_{req}})+\tau_{2}\frac{f_{FLOPs}(L_{c})}{w_{FLOPs}}-R,$ (2)
where the vector $\mathbf{d}=(d_{1},d_{2},\dots,d_{N})$ represents the data contribution of each client from its local dataset $\mathcal{D}_{n}$, while $d_{req}$ denotes the required amount of aggregated data contribution necessary to achieve an acceptable accuracy, which is a service-dependent requirement. $\tau_{1},\tau_{2}\geq 0$ denotes the weighting factor for the first and second terms, respectively. Moreover, $w_{FLOPs}$ is the computation workload to proceed with the full model $w$ during its forward and backward propagation for a single data sample, measured in Floating Point Operations Per Second (FLOPS). Here, $f_{FLOPs}(L_{c})$ as a function of the computation workload, measured in FLOPS, of the client-side model $w_{c}$ with respect to $L_{c}$. In this context, $f_{FLOPs}(L_{c})$ is the increasing function of $L_{c}$ since the computation workload increases with the number of layers. Nevertheless, due to the difficulty of analytically understanding this model-dependency, a data-driven approach can be applied to model its relationship through offline training using empirical measurements. Hence, we adopted regression-based modeling, which is one of the most widely used approaches, exemplified as mobile CPU properties, such as CPU power and temperature variation modeling [11, 12]. Consequently, as shown in Fig. 2, we conclude that $f_{FLOPs}(L_{c})$ is a linear function of $L_{c}$, modeled as $f_{FLOPs}(L_{c})=aL_{c}+b$, where $a$ and $b$ are coefficient values [^2] We used the thop package to measure the forward FLOPs and the model Params: https://pypi.org/project/thop/. Backword FLOPs is assumed to be twice the forward.. Finally, this utility function quantifies the balance between the satisfaction derived from aggregated data contribution from clients, and the gain from distributing training workload to clients, against the expenditure $R$ allocated to clients during the SFL process. The range of $R$ is set by the SFL model owner within $(0,R^{\max}]$, where $R^{\max}$ represents the upper limit of incentives regulated by the SFL model owner’s policy. Note that in the right-hand side of (2), the first term $\ln(1+\frac{\sum_{n=1}^{N}d_{n}}{d_{req}})$ suggests that the satisfaction gained from aggregated data contribution may follow the law of diminishing returns. From [13], the amount of raw data plays an important role in the quality of data analytics (e.g., accuracy). Moreover, numerous studies demonstrate that model accuracy increases concavely with the overall dataset size [14, 15, 16]. Accordingly, a concave, strictly increasing, and continuously differentiable function can represent this term. In this paper, we adopt a logarithmic utility function for this purpose, as it is widely used in the literature to quantify user satisfaction with diminishing returns.
II-B Utility Function of Clients
@@
II-B1 Overall Computation Energy of client
@@
II-B2 Overall Communication Energy of Client
To streamline our analysis and focus on the core intuition, we adopt a simplified approach similar to [17]. This involves assuming fixed uplink and downlink channels, maintaining a quasi-static channel mode throughout the SFL process facilitating efficient analysis. Then, the transmission latency, both i) for transmitting the smashed data from client $n$ to the main server and ii) for receiving the gradient data from the main server to the client $n$, is given by:
@@ II-B2 Overall Communication Energy of Client
Here, $s_{n}$ and $g_{n}$ denote the size of the smashed data and the gradient data of client $n$, respectively [^3] Here, we assume that the size of the smashed data is constant, although it could be model-dependent. For the evaluation, we set the size of the smashed data by averaging the size across layers to simplify the analysis.. Additionally, $r_{s:n}$ and $r_{g:n}$ represent the uplink transmission rate from client $n$ to the main server and the downlink transmission rate from the main server to client $n$, respectively. Similarly, the transmission latency, both i) for transmitting the client-side model from client $n$ to the fed server and ii) for receiving the global client-side model from the fed server to the client $n$, is given by:
@@
Here, $r_{cm:n}$ and $r_{gcm:n}$ represent the uplink transmission rate from client $n$ to the fed server and the downlink transmission rate from the fed server to client $n$, respectively. $m$ represents the number of bits for a single model parameter, where $|w_{c}|$ is the total number of model parameters in the client-side model. Note that the size of the global client-side model is the same as that of the individual client-side model. In this context, $|w_{c}|$ is the function of $L_{c}$ since the number of model parameters increases with the number of layers. Similar to $f_{FLOPs}$, due to the difficulty of analytically understanding, we also adopted regression-based modeling. Consequently, as shown in Fig. 2, we conclude that $|w_{c}|$ is a convex function of $L_{c}$, modeled as $|w_{c}|=c\exp{(dL_{c})}$, where $c$ and $d$ are coefficient values. Then, with given $p^{t}_{n}$ and $p^{r}_{n}$ of transmission and receiving power consumption of client $n$, overall communication energy for interacting with main server $E_{com-m:n}$ and fed server $E_{com-f:n}$ can be calculated by
@@
II-B3 Overall energy consumption
The energy consumption for $E$ epochs of client $n$ with $M$ mini-batches is given by $E_{n}^{epoch}=EM(E_{cmp,n}+E_{com-m,n})$, $n\in\mathcal{N}$. Consequently, the overall energy consumption for $T$ rounds, covering the entire duration until the SFL system completes its operations, is calculated by
@@ II-B3 Overall energy consumption
$\footnotesize E_{tot,n}=T((d_{n}f_{FLOPs}(L_{c})A_{n}+B_{n})+c\exp{(dL_{c})}C_{n}),n\in\mathcal{N},\vspace{-5pt}$ (9)
@@
$\footnotesize E_{tot,n}=d_{n}H_{n}+I_{n},n\in\mathcal{N},\vspace{-5pt}$ (10)
@@
II-B4 Shared Incentives and Utility Function Design
Adhering to the principle of proportional sharing, the SFL model owner allocates the amount of incentives $R$ to each client $n\in\mathcal{N}$ in proportion to their data contributions. Accordingly, the specific amount of incentives $R_{n}$ allocated to client $n$ by the SFL model owner is determined as
@@ II-B4 Shared Incentives and Utility Function Design
Thus, the utility function of the client $n$ is defined as
@@
$=\psi_{n}R\frac{d_{n}}{\sum_{l=1}^{N}d_{l}}-d_{n}H_{n}-I_{n}+S,$ (12)
where $\psi_{n}(>0)$ is the weighting factor for the incentives of client $n$. And $S$ is an offset to $U_{n}$ as a baseline constant term, representing the satisfaction of clients using the SFL model.
Figure 3: A diagram of two-stage Stackelberg game in SFL where SFL model owner is a leader and clients are followers.
III Game-Theoretic Analysis
In this section, we examine the competitive dynamics between the SFL model owner and the clients, while also considering the competition among clients to gain more incentives. These competitive behaviors can be considered as a hierarchical noncooperative decision problem, analyzable as a Stackelberg game. Stackelberg games are a type of noncooperative game characterized by a hierarchy among players, where they are classified as leaders and followers [18, 19, 20]. In this study, we consider a scenario where the clients follow the behavior of the SFL model owner. Treating the SFL model owner as the leader and the clients as the followers, the competition among players to maximize their utility is modeled as a two-level Stackelberg game. If the SFL model owner decides on the cut layer and the total amount of incentives, the clients will respond by selecting their optimal strategies regarding data contribution for the SFL process. The SFL model owner is aware that the clients will choose their best response strategies based on the owner’s decisions. Therefore, the SFL model owner optimizes its strategy to maximize utility, taking into account the clients’ best responses. To find the Stackelberg Equilibrium (SE), which is the solution to this game, we employ a backward induction technique introduced in [18, 19]. As illustrated in Fig. 3, this technique can be summarized as follows. First, in the lower-level game for the clients, referred to as Stage 1: Data Contribution Management Competition, the clients determine the amount of data to contribute to maximize their utility function (II-B4), given the incentives and cut layer configuration. Then, in the upper-level game for the SFL model owner, referred to as Stage 2: Maximizing Utility of SFL Model Owner, the SFL model owner combines the clients’ strategies with its utility function (2) and selects the cut layer and incentive amounts that maximize its own utility function (2).
III-A Stage 1: Noncooperative Game Among Clients
Each client aims to maximize the utility function (II-B4) within the dataset constraints [0, $D_{n}$]. With given the vector $\mathbf{d}=(d_{1},d_{2},\dots,d_{N})$, let $\mathbf{d}_{-n}$ denote the profile of strategies for the clients except $n$, i.e., $\mathbf{d}_{-n}=(d_{l})_{l\in\mathcal{N}\backslash\{n\}}$. Then, by definition, we have $\mathbf{d}=(d_{n},\mathbf{d}_{-n})$.

### v3

# How Can Incentives and Cut Layer Selection Influence Data Contribution in Split Federated Learning?

@@ Abstract
Split Federated Learning, Wireless Networking, Game Theory, Incentive Mechanism, Cut Layer Selection.
@@ I Introduction
  • FL: FL allows a server model to be trained in parallel across many clients, but each client is required to run an entire model, resulting in a significant training burden. This leads to high battery consumption on clients, significant client-side computation requirements, and limited model privacy.
@@
  • Our findings show that the Nash and Stackelberg strategies are desirable operating points for all participants (i.e., the SFL model owner and clients) in competitive situations. From our classification experiments using the CIFAR-10 dataset, we highlight the importance of incentive mechanisms in accelerating SFL and enhancing model accuracy. In our experiment, accuracy varies from 60% to 91% depending on the incentives provided. Specifically, we evaluate the effectiveness of these strategies under various parameter settings, which motivates the design of incentive mechanisms for practical SFL applications.
The remainder of this paper is organized as follows. Section II reviews related work on cut layer selection and resource management for SFL. Section III presents the system model and formulates the utility functions for both the SFL model owner and the clients. In Section IV, we describe the competitive game between the SFL model owner and the clients, rigorously formulating the problem as a Stackelberg game. We then characterize the closed-form solutions for a Nash Equilibrium of the clients and provide the algorithm to obtain a Stackelberg Equilibrium. Section V provides the performance evaluation. Section VI discusses the impact of cut layer selection and differential privacy on the privacy level, offering guidelines for setting the minimum cut layer requirements during the SFL process. Finally, Section VII concludes the paper and suggests directions for future research.
@@
II Related Work
To mitigate the training burden in FL and improve convergence speed, SFL has emerged as a promising method, integrating the benefits of federated and split learning. However, SFL is associated with substantial communication overhead, mainly due to the transfer of smashed data, gradients, and model updates between clients, the main server, and the federated server. This necessitates enhancements [7]. Recent studies have focused on the impact of cut layer selection on system performance and privacy. Notably, [5, 2] developed analytical models to explore the effect of total model training time (latency) on the cut layer point. Based on these analyses, various strategies have been proposed to enhance communication efficiency in SFL [7, 5, 6]. For instance, [5] introduced an optimal cut layer selection method based on latency analysis, followed by a convergence analysis of the SFL framework. Extending this work, [6] examined the latency of SFL in wireless networks and designed a joint optimization problem involving cut layer selection, client clustering, and bandwidth allocation to minimize overall training latency. Similarly, [10] proposed a joint method for cut layer selection and server computation resource allocation for clients. [7] employed an auxiliary network for local updates of client-side models, maintaining only a single server-side model to reduce storage costs on the main server, and conducted a comprehensive analysis of communication costs, storage space, and convergence. Additionally, [8] focused on reducing the dimensions of activations’ gradients for backpropagation to decrease communication overhead. This study also explored joint optimization strategies for subchannel allocation, power control, and cut layer selection in wireless networks, aiming to minimize per-round latency. The recent work of [11] utilized SFL for Unmanned Aerial Vehicles (UAVs) networks considering wireless network perspectives. It generated the mobile traffic prediction model from UAVs collaboratively by selecting the cut layer in SFL for energy-efficient training.
In addition to these efficiency considerations, SFL poses privacy concerns due to interactions between clients and servers, including both the main and federated servers. During the forward phase, clients transmit smashed data to the main server, which becomes vulnerable to reconstruction attacks, potentially compromising the privacy of the original information. Furthermore, exchanging model updates with the federated server introduces additional risks to raw data privacy. To address these issues, several privacy-preserving mechanisms have been proposed. For instance, a recent study [12] on split learning empirically investigated the impact of cut layer selection on reconstruction attacks during the forward phase. The findings suggested that having deeper layers on the client side, which involve more non-linear functions, compresses raw data by eliminating less informative features, thereby enhancing resistance to reconstruction attacks. Similarly, [13] conducted empirical studies on how the distance correlation between raw and smashed data affects privacy leakage in split learning, emphasizing the importance of cut layer selection. Recently, the work in [14] analyzed the trade-off between energy consumption and privacy levels in relation to cut-layer selection in SFL. Furthermore, the work of [15] introduced an innovative cut layer selection scheme aimed at reducing the overall training latency in SFL, taking into account wireless network perspectives. Specifically, the proposed scheme not only minimized the overall training latency but also ensured an acceptable level of privacy by setting boundary constraints on the cut layer selection.
However, none of these studies address the strategic mechanisms of the SFL model owner and clients considering their competitive behaviors, which is one of our contributions. Notably, our approach to formulating and analyzing a Stackelberg game with two solution concepts involving the SFL model owner and clients is not fully covered in previous work.
Figure 2: Timeline of the Overall SFL Process.
III System Model
@@
The SFL framework consists of two servers (which can be either physical or logical servers): (i) a federated learning server, referred to as the fed server, and (ii) a main server, along with multiple clients participating as part of the SFL process. In this setup, we define a set of clients as $\mathcal{N}={1,2,\cdots,N}$, where each client is represented as $n\in\mathcal{N}$ and has its own local dataset $\mathcal{D}_{n}$. The size of each dataset $\mathcal{D}_{n}$ is indicated by the number of data items, denoted as $D_{n}=|\mathcal{D}_{n}|$, with each data item comprising both features and labels. The entire model, denoted as $w$, is composed of $L$ layers. Specifically, the model $w$ is divided into two sub-models at the $L_{c}$-th layer, referred to as the cut layer. These sub-models are: (i) the client-side model $w_{c}$, including layers from the first layer up to layer $L_{c}$, and (ii) the server-side model $w_{s}$, encompassing layers from layer $L_{c}+1$ to layer $L$. Thus, the full model can be represented as $w=[w_{c};w_{s}]$. As shown in Fig. 1, each client $n$ in the set $\mathcal{N}$ maintains its client-side model, denoted as $w_{c,n}$. These client-side models are trained in parallel using the clients’ respective local datasets, through interactions with the server-side model $w_{s}$ hosted on the main server. To synchronize the client-side models, clients participate in federated learning, coordinated by the fed server, to construct a global client-side model $w_{c}$ in a collaborative manner [16].
@@
Finally, the entire SFL process is carried out over $T$ rounds. At the beginning of each round $t$, clients are provided with the latest global client-side model $w_{c}$ from the fed server. Following this, both the clients and the main server engage in training over $E$ epochs, updating their respective client-side models $w_{c,n}$ for each client $n\in\mathcal{N}$, and the server-side model $w_{s}$. Then, each client $n$ randomly samples a mini-batch instance $\zeta_{n}$ from its dataset $\mathcal{D}_{n}$. Note that during each epoch, clients randomly partition their datasets into $M$ mini-batches. Stochastic gradient descent (SGD) is applied to each mini-batch, allowing clients to update their client-side models $w_{c,n}$ and the main server to update the server-side model $w_{s}$ in a coordinated fashion. Once the $E$ epochs are concluded, the fed server collects the updated client-side models $w_{c,n}$ from all clients and combines them to form an updated global client-side model $w_{c}$ [^1] At mini-batch instance $\zeta_{n}$, $l_{n}(w;\zeta_{n})$ denotes the loss of the model $w$ over client $n$’s mini-batch instance $\zeta_{n}$, which can be expressed as follows: $l_{n}(w;\zeta_{n})=h(w_{s};u(w_{c,n};\zeta_{n})).$ (1) Here, $u$ is a function on the client side that transforms the input data from the sampled mini-batch instance $\zeta_{n}$ into the activation space, producing the output of the client-side model $w_{c,n}$ (i.e., smashed data). Meanwhile, $h$ is a function on the server side that takes the activation space as input and produces a single scalar value representing the loss.. A detailed timeline of overall SFL process is summarized in Fig. 2.
III-A Utility Function of SFL Model Owner
@@
$U_{MO}(R,L_{c},\textbf{d})=\tau_{1}\ln\left(1+\frac{\sum_{n=1}^{N}d_{n}}{d_{req}}\right)+\tau_{2}\frac{f_{FLOPs}(L_{c})}{w_{FLOPs}}-R,$ (2)
where the vector $\mathbf{d}=(d_{1},d_{2},\dots,d_{N})$ contains the data contribution of each client from its local dataset $\mathcal{D}_{n}$, while $d_{req}$ denotes the required amount of aggregated data contribution necessary to achieve an acceptable accuracy, which is a service-dependent requirement. $\tau_{1},\tau_{2}\geq 0$ are weights for trading of the first and second terms. Moreover, $w_{FLOPs}$ is the computation workload to proceed with the full model $w$ during its forward and backward propagation for a single data sample, measured in Floating Point Operations Per Second (FLOPS). Here, $f_{FLOPs}(L_{c})$ as a function of the computation workload, measured in FLOPS, of the client-side model $w_{c}$ with respect to $L_{c}$. In this context, $f_{FLOPs}(L_{c})$ is an increasing function of $L_{c}$ since the computation workload increases with the number of layers. Nevertheless, due to the difficulty of analytically characterizing this model-dependency, a data-driven approach can be applied to model its relationship through offline training using empirical measurements. Hence, we adopt regression-based modeling, which is one of the most widely used approaches for modeling mobile CPU properties, such as CPU power and temperature variation modeling [17, 18]. Consequently, as shown in Fig. 3, we conclude that $f_{FLOPs}(L_{c})$ is an affine function of $L_{c}$, modeled as $f_{FLOPs}(L_{c})=aL_{c}+b$, where $a$ and $b$ are coefficient values [^2] We used the thop package to measure the forward FLOPs and the model Params: https://pypi.org/project/thop/. Backword FLOPs are assumed to be twice the forward FLOPs.. Finally, this utility function quantifies the balance between the satisfaction derived from aggregated data contribution from clients, and the gain from distributing training workload to clients, against the expenditure $R$ allocated to clients during the SFL process. The range of $R$ is set by the SFL model owner within $(0,R_{\max}]$, where $R_{\max}$ represents the upper limit of incentives regulated by the SFL model owner’s policy. Note that in the right-hand side of (2), the first term $\ln(1+\frac{\sum_{n=1}^{N}d_{n}}{d_{req}})$ suggests that the satisfaction gained from aggregated data contribution may follow the law of diminishing returns. From [19], the amount of raw data plays an important role in the quality of data analytics (e.g., accuracy). Moreover, numerous studies demonstrate that model accuracy increases concavely with the overall dataset size [20, 21, 22]. Accordingly, a concave, strictly increasing, and continuously differentiable function can represent this term. In this paper, we adopt a logarithmic utility function for this purpose, as it is widely used in the literature to quantify user satisfaction with diminishing returns.
III-B Utility Function of Clients
@@
III-B1 Overall computation energy of the client
@@
III-B2 Overall communication energy of the client
To streamline our analysis and focus on the core intuition, we adopt a simplified approach similar to [23]. This involves assuming fixed uplink and downlink channels, maintaining a quasi-static channel mode throughout the SFL process facilitating efficient analysis. Then, transmission latencies for transmitting the smashed data from client $n$ to the main server and for receiving the gradient data from the main server to the client $n$, are given by:
@@ III-B2 Overall communication energy of the client
Here, $s_{n}$ and $g_{n}$ denote the size of the smashed data and the gradient data of client $n$, respectively [^3] Here, we assume that the size of the smashed data is constant, although it could be model-dependent. For the evaluation, we set the size of the smashed data by averaging the size across layers to simplify the analysis.. Additionally, $r_{s:n}$ and $r_{g:n}$ represent the uplink transmission rate from client $n$ to the main server and the downlink transmission rate from the main server to client $n$, respectively. Similarly, transmission latencies for transmitting the client-side model from client $n$ to the fed server and for receiving the global client-side model from the fed server to the client $n$, are given by:
@@
Here, $r_{cm:n}$ and $r_{gcm:n}$ represent the uplink transmission rate from client $n$ to the fed server and the downlink transmission rate from the fed server to client $n$, respectively. $m$ represents the number of bits for a single model parameter, where $|w_{c}|$ is the total number of model parameters in the client-side model. Note that the size of the global client-side model is the same as that of the individual client-side model. In this context, $|w_{c}|$ is the function of $L_{c}$ since the number of model parameters increases with the number of layers. Similar to $f_{FLOPs}$, due to the difficulty of analytical modeling, we also adopted regression-based modeling. Consequently, as shown in Fig. 3, we conclude that $|w_{c}|$ is a convex function of $L_{c}$, modeled as $|w_{c}|=c\exp{(dL_{c})}$, where $c$ and $d$ are coefficient values. Then, with given $p^{t}_{n}$ and $p^{r}_{n}$ of transmission and receiving power consumption of client $n$, the overall communication energy for interacting with main server $E_{com-m:n}$ and fed server $E_{com-f:n}$ can be calculated by
@@
III-B3 Overall energy consumption
The energy consumption for $E$ epochs of client $n$ with $M$ mini-batches is given by $E_{n}^{epoch}=EM(E_{cmp,n}+E_{com-m,n})$, $n\in\mathcal{N}$. Consequently, the overall energy consumption for $T$ rounds, covering the entire duration until the SFL system completes its operations, is given by
@@ III-B3 Overall energy consumption
$\footnotesize E_{tot,n}=T((d_{n}f_{FLOPs}(L_{c})A_{n}+B_{n})+c\exp{(dL_{c})}C_{n}),n\in\mathcal{N},\vspace{-1pt}$ (9)
@@
$\footnotesize E_{tot,n}=d_{n}H_{n}+I_{n},n\in\mathcal{N},\vspace{-1pt}$ (10)
@@
III-B4 Shared incentives and utility function design
Adhering to the principle of proportional sharing, the SFL model owner allocates an amount of incentives $R$ to each client $n\in\mathcal{N}$ in proportion to their data contributions. Accordingly, the specific amount of incentives $R_{n}$ allocated to client $n$ by the SFL model owner is determined as
@@ III-B4 Shared incentives and utility function design
Thus, the utility function of client $n$ is given by
@@
, $=\psi_{n}R\frac{d_{n}}{\sum_{l=1}^{N}d_{l}}-d_{n}H_{n}-I_{n}+S,$ (12)
where $\psi_{n}(>0)$ is a weighting factor for the incentives of client $n$. And $S$ is an offset to $U_{n}$ as a baseline constant term, representing the satisfaction of clients using the SFL model.
Figure 4: A diagram of a two-stage Stackelberg game in SFL where the SFL model owner is a leader and clients are followers.
IV Game-Theoretic Analysis
In this section, we examine the competitive dynamics between the SFL model owner and the clients, while also considering the competition among clients to gain more incentives. These competitive behaviors can be considered as a hierarchical noncooperative decision problem, analyzable as a Stackelberg game. Stackelberg games are a type of noncooperative game characterized by a hierarchy among players, in which they are classified as leaders and followers [24, 25, 26]. In this study, we consider a scenario where the clients follow the behavior of the SFL model owner. Treating the SFL model owner as the leader and the clients as the followers, the competition among players to maximize their utilities is modeled as a two-level Stackelberg game. If the SFL model owner decides on the cut layer and the total amount of incentives, the clients will respond by selecting their optimal strategies regarding data contribution for the SFL process. The SFL model owner is aware that the clients will choose their best response strategies based on the owner’s decisions. Therefore, the SFL model owner optimizes its strategy to maximize utility, taking into account the clients’ best responses. To find the Stackelberg Equilibrium (SE), which is the solution to this game, we employ a backward induction technique introduced in [24, 25]. As illustrated in Fig. 4, this technique can be summarized as follows. First, in the lower-level game for the clients, referred to as Stage 1: Data Contribution Management Competition, the clients determine the amount of data to contribute to maximize their utility function (III-B4), given the incentives and cut layer configuration. Then, in the upper-level game for the SFL model owner, referred to as Stage 2: Maximizing Utility of SFL Model Owner, the SFL model owner combines the clients’ strategies with its utility function (2) and selects the cut layer and incentive amounts that maximize its own utility function (2).
IV-A Stage 1: Noncooperative Game Among Clients
Each client aims to maximize its utility function (III-B4) within the dataset constraints [0, $D_{n}$]. Given the vector $\mathbf{d}=(d_{1},d_{2},\dots,d_{N})$, let $\mathbf{d}_{-n}$ denote the profile of strategies for the clients except $n$, i.e., $\mathbf{d}_{-n}=(d_{l})_{l\in\mathcal{N}\backslash\{n\}}$. Then, by definition, we have $\mathbf{d}=(d_{n},\mathbf{d}_{-n})$.